OpenStudy (anonymous):

San Carlos is 52 miles north of Cross Creek. Cottonwood is 75 miles west of Cross Creek. How far apart are San Carlos and Cottonwood? Round your answer to the nearest mile. A. 33 miles B. 91 miles C. 106 miles D. 127 miles @OrangeMaster

OpenStudy (anonymous):

A? @OrangeMaster

OpenStudy (anonymous):

For Q4 remember that N-S and E-W lines cross over at right angles, put a dot on a piece of paper representing Cross Creek then dots for the other two locations.

OpenStudy (anonymous):

I got 91.

OpenStudy (anonymous):

Do you want to see the math @CrazyCountryGirl

OpenStudy (anonymous):

52*52=2704 75*75=5625 5625+2704=8329 \[\sqrt{8329}=91\]

OpenStudy (anonymous):

Um...Yes please. :)

OpenStudy (anonymous):

I am not 100% sure, but I think that C should always be longer than A and B

OpenStudy (anonymous):

So the answer A should be ruled out from the beginning.

OpenStudy (anonymous):

What is the length of the diagonal of a rectangle whose width is 100 centimeters and whose length is 180 centimeters? (Round your answer to the nearest whole centimeter.) A. 202 cm B. 206 cm C. 209 cm D. 212 cm A???????????????????????????

OpenStudy (anonymous):

Close, it's B. I'll show the math is a sec.

OpenStudy (anonymous):

Yeah, I though it was A or B. :)

OpenStudy (anonymous):

100*100=10,000 180*180=32,400 32,400+10,000=42,400\[\sqrt{42,400}=205.9\]205.9 gets rounded to 206

OpenStudy (anonymous):

Sorry, I got the answers. I thought A was 206. :) I meant B, I was like I was so sure I was right, and so I looked at the answer choices and I was like ooooohhhhh that makes sense now. :)

OpenStudy (anonymous):

Sorry, I got the answer mixed up

OpenStudy (anonymous):

OK, in that case then. Good job.

OpenStudy (anonymous):

yay! thx!!

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

Do you think that you feel comfortable using the Pythagorean Theorem now.

OpenStudy (anonymous):

I got a 100! Yeah, thanks. :)

OpenStudy (anonymous):

No problem, glad I could teach another brain.

OpenStudy (anonymous):

Once you get the hang of it, it should be super easy and fast.